A survey of 98 randomly selected students finds that they spend a mean of $435 per semester. Assume that the population standard deviation is $56 per month.

Confidence Interval: What is the 95% confidence interval to estimate the population mean? (Round your answers to two decimal places.)

______ < u < _______


1
Expert's answer
2022-05-10T23:46:08-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\dfrac{\sigma}{\sqrt{n}})"

"=(435-1.96\dfrac{56}{\sqrt{98}},435+1.96\dfrac{56}{\sqrt{98}})"

"=(423.913, 446.087)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "423.913 < \mu < 446.087," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(423.913, 446.087)."